OpenStudy (lilkg77):

Which of the following recursive formulas represents the same geometric sequence as the formula an = 2 • 2n - 1? A. an = 2 • an - 1 B. an = 4 • an - 1 C. an = 4 + an - 1 D. an = 2 + an - 1

OpenStudy (sparrow2):

it's b

OpenStudy (sparrow2):

wait,what is a?

OpenStudy (lilkg77):

@sparrow2 what do u mean?

OpenStudy (sparrow2):

is it a*n=2*2*n-1?

OpenStudy (lilkg77):

The n is below the a, idk what its called

OpenStudy (lilkg77):

@sparrow2

OpenStudy (sparrow2):

index

OpenStudy (sparrow2):

are you sure that in B there is an not only n?

OpenStudy (xapproachesinfinity):

hey lil the geometric you are given is it like this \(a_n=2~2^{n-1}\)

OpenStudy (lilkg77):

Yes @xapproachesinfinity exactly like that

OpenStudy (xapproachesinfinity):

ok good

OpenStudy (xapproachesinfinity):

to know which recursive formula work we need to try some values and compare

OpenStudy (xapproachesinfinity):

or try to come up with the recursive formula for the given implicit formula

OpenStudy (xapproachesinfinity):

let's try the latter one. we have \[a_n=2~~2^{n-1}\] so \(a_{n+1}=2~2^{n}\)

OpenStudy (xapproachesinfinity):

hey where did you go